OpenStudy (anonymous):

suppose that an individuals starting salary is $35000 per year and that individual recieves a $2000 raise each year. Find the total amount earned over 20 years.

OpenStudy (anonymous):

35000 + 2000(k+1) from 0 - 20..something like that series

OpenStudy (anonymous):

20*35,000+1(2000)+2(2000)+3(2000)+.....+20(2000)

OpenStudy (anonymous):

S20=10(2(35000)+(19)2000)=$1080000

OpenStudy (anonymous):

(35,000)+(35000+2000)+(35,000+4000)+(35000+6000)+......(35000+38000) =35000*20+(190*2000)=1080000 * I got 190 by adding 1 to 19

OpenStudy (anonymous):

thank u...